A 11.38 L balloon contains 0.63 mol of air at 185.50 kPa pressure. What is the temperature of the air in the balloon?



1
Expert's answer
2022-04-07T04:06:00-0400

P1V1/T1 = P2V2/T2

P1*6.3/T1 = 1.5P1*V2/0.75T1

6.3 = 2V2

V2 = 6.3 / 2 = 3.15 (L)